OpenStudy (anonymous):

the sum of 2 numbers is 74. The larger number is 3 more than the smaller. find the numbers

OpenStudy (anonymous):

x + (x + 3) = 74 2x + 3 = 74 2x = 71 x = 35.5 35.5 + 3 = 38.5 The two numbers are 35.5 and 38.5

OpenStudy (anonymous):

yes i know but im supposed to solve the equation with 2 variables

OpenStudy (anonymous):

Ok, how about x + y = 74 y = x + 3 Then just substitute the x + 3 in for y.

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

Basically, you use substitution. Instead of x + 3, you use y. Then, put in another equation that solves for y which is given to you. THis forms a system of linear equations. It's just adding in another system of equations.